You are here: Foswiki>FLAO Web>LbtTel>TelescopeLogLBT672a>TelLogLBT672a20100511 (12 May 2010, FernandoQuiros)Edit Attach

Interferometer alignment

G. Brusa, R. Briguglio, L. Fini, F. Quiros.

Starting time: 7am.

System initialization problem:
  • The hexapod wakes up in an wierd state (OSS GUI displayed hexapod position NOT EQUAL TO actual hexapod position):
    • Probably, there was a short hexapod power loss suffered in the night (Runa investigating the issue looking at the logs).
  • The OSS of the Secondary mirror says: state: In position, status: No Error. However, the hexapod position vector [x,y,z,rx,ry,rz] is all zeros!!!!!!!!! On the other hand, we know the actual hexapod position is the one in which the hexapod was just before the power off (confirmed by the fact that we see light arriving to the interferometer as yesterday).
  • Some action items arise from this problem:
    • OSS GUI: The message state: In position, status: No Error with all zeros is not correct! Luca in charge of sending an email to Paul to report this.

9am: GENERAL POWER SHUTDOWN IN THE OBSERVATORY SCHEDULED TO SWAP FROM COMMERCIAL POWER SUPPLY TO THE OBSERVATORY GENERATOR........

After power on we continued to investigate the problem of the hexapod position:
  • PROBLEM DETECTED: The secondary mirror was left in simulation mode at the end of the night observations!!!!! grrrrrrr......

12pm: PAUSE TO GIVE A TOUR OF THE TELESCOPE TO VISITORS FROM 4D, and lunch.

After lunch we try proceed (or better said "start") with our work.

2pm: We continue with technical problems:
  • ADAM power off: The ADAM continues to power off and powering on itself after about 0.5 seconds (this has happened several times already at least from yesterday).
  • 3pm: Actuator 112 is going mad: the shell is ripped...
  • 3.23pm: Actuator 75 had a current error.

PRELIMINARY ALIGNMENT SUCCESSFUL:

Finally, in the rest of the day we managed to improve substantially the alignment of the interferometer by moving both the hexapod and the tertiary. The procedure followed to achieve this was based on the calibration of an interaction matrix between the degrees of freedom and the first Zernike modes (tip, tilt, focus, and two comas) measured with the interferometer. Proceeding in small steps, with the use of the corresponding reconstruction matrix, we managed to reduce the residual Zernike coefficients sufficiently enough to start taking some measurements of the mirror modes...

PRELIMINARY MIRROR MODES MEASUREMENTS:

After dinner, and taking advantage that the wind speed is at its 'normal' levels (that is 30 m/s), we were allowed to continue working for disgrace of our fellow astronomers. Runa starts to take some sets of measurements, and believe it or not some technical problems arise:
  • 8:43pm: Laser illumination is oscillating... about 20 cycles.
  • measurements acquisiton bug: some measurements get lost. Runa looking at it.
-- FernandoQuiros - 11 May 2010
Topic revision: r5 - 12 May 2010, FernandoQuiros
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback